OpenStudy (anonymous):

Find the indicated values where g(t)=t^2-t and f(x)=1+x g(f(0))+f(g(0))

OpenStudy (shamil98):

g(f(0)) = evaluate f(0) and then put the result into the equation for g(t). f(g(0)) evaluate g(0) and put into the equation for f(x) then add those two together

OpenStudy (anonymous):

What is f(0)? @shamil98

OpenStudy (shamil98):

f(x) = 1 + x f(0) = 1 + 0 = 1 you substitute the value for the x's in the equation and simplify.

OpenStudy (shamil98):

same thing for g(0)

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

@shamil98 So would the answer be 1?

OpenStudy (shamil98):

yeah

OpenStudy (shamil98):

g(f(0))+f(g(0)) = 0 + 1 = 1

OpenStudy (anonymous):

Thank you!